Theory and Applications of Categories, Vol. 30, No. 7, 2015, pp. 206–228.

CHARACTERISTIC SUBOBJECTS IN SEMI-ABELIAN CATEGORIES

ALAN S. CIGOLI AND ANDREA MONTOLI

Abstract. We extend to semi-abelian categories the notion of characteristic subobject, which is widely used in group theory and in the theory of Lie algebras. Moreover, we show that many of the classical properties of characteristic subgroups of a group hold in the general semi-abelian context, or in stronger ones.

1. Introduction

The notion of characteristic subgroup (which means a subgroup that is invariant under all automorphisms of the bigger group) is widely used in group theory. Examples of characteristic subgroups are the centre and the derived subgroup of any group. The main properties of characteristic subgroups are the following: if H is a characteristic subgroup of K and K is a characteristic subgroup of G, then H is a characteristic subgroup of G; moreover, if H is characteristic in K and K is normal in G, then H is normal in G. These transitivity properties of characteristic subgroups imply, for example, that the derived series and the central series of a group are normal series, and this fact is very useful in order to deal with solvable and nilpotent groups.

An analogous notion exists for Lie algebras (over a commutative ring R): a characteristic ideal of a Lie algebra is a subalgebra which is invariant under all derivations of the bigger one. The two transitivity properties mentioned above hold also in this context, and again this allows to easily describe solvable and nilpotent Lie algebras.

The strong parallelism between these two contexts is explained by the fact that automorphisms represent group actions, while derivations represent actions of Lie algebras in the following sense. An action of a group B on a group G can be described simply as a group homomorphism B → Aut(G); in the same way, an action of a Lie algebra B on a Lie algebra G is a homomorphism of Lie algebras B → Der(G).

The first author was partially supported by FSE, Regione Lombardia. The second author was partially supported by the Centro de Matemática da Universidade de Coimbra (CMUC), funded by the European Regional Development Fund through the program COMPETE and by the Portuguese Government through the FCT - Fundação para a Ciência e a Tecnologia under the projects PEst-C/MAT/UI0324/2013 and PTDC/MAT/120222/2010 and grant number SFRH/BPD/69661/2010.

Received by the editors 2014-02-12 and, in revised form, 2015-02-09.

Transmitted by Stephen Lack. Published on 2015-02-10.

2010 Mathematics Subject Classification: 17B30, 17D99, 18C05, 08A30, 08C05.

Key words and phrases: characteristic subobject, semi-abelian categories, commutators, centralisers.

\begin{matrix}c \\ ◯ \end{matrix} Alan S. Cigoli and Andrea Montoli, 2015. Permission to copy for private use granted.

The aim of this paper is to extend the definition and the main properties of characteristic subobjects to the context of semi-abelian categories [18]. Examples of semi-abelian categories are groups, rings, associative algebras, Lie algebras and, in general, any variety of Ω-groups. Our definition is based on the notion of internal action introduced in [3]. In [9] it is proved that, in semi-abelian categories, internal actions are equivalent to split extensions, via a semidirect product construction which generalises the classical one known for groups.

We define a characteristic subobject as a subobject H of an object G which is invariant under all (internal) actions on G. In the semi-abelian context, we can use the equivalence between actions and split extensions mentioned above in order to deduce properties of characteristic subobjects from properties of the kernel functor which associates with any split epimorphism its kernel.

The paper is organized as follows: in Section 2 we give the definition of characteristic subobject and we prove some properties that hold in any semi-abelian category, like the transitivity properties mentioned at the beginning, or the fact that the intersection and the join of two characteristic subobjects is characteristic. Then we study properties that hold in stronger contexts, such as:

- the commutator of two characteristic subobjects is characteristic (Section 3);

- the centraliser of a characteristic subobject is characteristic (Section 4).

Some properties about actors of characteristic subobjects are studied in Section 5 in the context of action representative categories [4, 2] and analogous results are proved in action accessible categories [10], replacing actors with suitable objects.

2. Definition and basic properties

A characteristic subgroup of a group G is classically defined as a subgroup H of G which is invariant under all the automorphisms of G. This means that any automorphism of G restricts to an automorphism of H. Since the automorphism group Aut(G) of a group G classifies all the group actions on G, a subgroup H of a group G is characteristic if and only if any group action on G restricts to an action on H.

In other algebraic contexts it is no longer true that automorphisms classify actions, hence the notions of invariance under automorphisms and invariance under actions are different. As already explained in the introduction, here we are interested in the latter. In order to study it in the semi-abelian setting, we are going to use the notion of internal action, introduced in [3]. Let us briefly recall the definition.

Let 𝒞 be a pointed category with finite limits and finite coproducts. For any object B in 𝒞, we can define the category Pt_{B}(𝒞) of points over B, whose objects are split epimorphisms (A, p, s) with codomain B and whose arrows are commutative triangles of the following form, with p′f = p and fs = s′:

\xymatrix{A \ar[rr]^{f} \ar[dr]^{p} &  & A′ \ar[dl]^{p′} \\  & B \ar[ul]^{s} \ar[ur]^{s′}}

We then get the two following functors:

Ker_{B}: Pt_{B}(𝒞) → 𝒞 ,

given by Ker_{B}(A, p, s) = Ker p, and

B + (−): 𝒞 → Pt_{B}(𝒞) ,

where B + (X) is the point \xymatrix{B + X \ar[r]^{[1,0]} & B \ar[l]^{ι_{B}}} .

These functors give rise to an adjunction. The corresponding monad on 𝒞 is denoted by B♭(−). For any object X ∈ 𝒞, we have that B♭X is the kernel of the morphism [1, 0]: B + X → B. The algebras for this monad are called internal B-actions. The comparison functor associates with every point (A, p, s) an action ξ as described in the following diagram (where X is the kernel of p and ξ is induced by the universal property of X):

[figure: B♭X; ker[1,0]; ξ; B + X; [1,0]; [s,k]; B; ι_{B}; X; k; A; p; B; s]

When 𝒞 is the category Gp of groups, the elements of B♭X are generated by formal sequences of type (b; x; b^{−1}) with b ∈ B and x ∈ X, and the internal action ξ is nothing but the realisation of these sequences in X, that is ξ(b; x; b^{−1}) = bxb^{−1}, or more properly ξ(b; x; b^{−1}) = k^{−1}(s(b)k(x)s(b^{−1})) since the product is actually computed in A.

Vice versa, given a group action ξ of B over X, we can always associate with it the semidirect product X ⋊_{ξ} B and a morphism of split extensions as in the following diagram:

[figure: B♭X; ker[1,0]; ξ; B + X; [1,0]; B; ι_{B}; X; i_{X}; X ⋊_{ξ} B; p_{B}; B; i_{B}]

We can repeat the same construction in any pointed category with finite limits, finite coproducts and, in addition, coequalisers, defining the object X ⋊_{ξ} B by means of the following coequaliser diagram:

\xymatrix{B♭X \ar[r]^{ker[1,0]} \ar[r]_{ι_{X}ξ} & B + X \ar@{->>}[r] & X ⋊_{ξ} B .}

However, in general, the sequence (i_{X}, p_{B}) above is not a short exact sequence. This is the case when the comparison functor between points and actions is an equivalence, i.e. when a categorical semidirect product in the sense of [9] is defined, as, for example, in any semi-abelian category.

We are now ready to give the following definition:

2.1. Definition. Let 𝒞 be a semi-abelian category, G an object in 𝒞 and h: H ↣ G a subobject. We say that H is characteristic in G, and we write H <_{char} G, if, for each pair

(B, ξ), with B an object of 𝒞 and ξ an internal action of B on G, the action ξ restricts to the subobject H. In other words, there exists a (unique) action \overline{ξ} of B on H which makes the following diagram commute:

[figure: B♭H; \overline{ξ}; 1♭h; B♭G; ξ; H; h; G]

Being 𝒞 a semi-abelian category, the above mentioned equivalence between actions and points allows us to reformulate the definition of characteristic subobject.

2.2. Proposition. Let 𝒞 be a semi-abelian category. A subobject h: H ↣ G is characteristic in G if and only if, for every split extension of kernel G

[figure: G; X; B] ,

there exist a split extension [figure: H; Y; B] and a morphism of split extensions as below, whose components on kernels and on cokernels are h and 1_{B} respectively (it is necessarily a monomorphism):

[figure: H; h; Y; B; G; X; B]  (1)

As we will see afterwards, this reformulation makes the notion of characteristic subobject much easier to handle. Moreover, the translation in terms of points reveals that, when actions are equivalent to points, many properties of characteristic subobjects are strictly related to the properties of the fibration of points (see [1]) or, to be more precise, of the kernel functors:

Ker_{B}: Pt_{B}(𝒞) → 𝒞 .

2.3. Proposition. If H is a characteristic subobject of K, and K is a characteristic subobject of G, then H is characteristic in G.

Proof. The result is a straightforward consequence of Definition 2.1.

2.4. Proposition. If H is a characteristic subobject of K, and K is a normal subobject of G, then H is normal in G.

Proof. It suffices to observe that, in the semi-abelian context, normal subobjects are exactly those closed under the conjugation action (i.e. clots, see for example [20]). Indeed, the conjugation action of G on itself restricts to K by normality, and then to H, since H <_{char} K, thus proving that H ◁ G.

2.5. Corollary. If H is a characteristic subobject of G, then H is normal in G.

2.6. Remark.

1. Combining Propositions 2.2 and 2.4, we obtain that every characteristic subobject h: H ↣ G gives rise, for every action of an object B on G, to a normal monomorphism of split extensions. Indeed, considering the morphism in diagram (1), we have that, since H <_{char} G, then H ◁ X; according to [1, Proposition 6.2.1], this suffices to

prove that (1) is a normal monomorphism in Pt_{B}(𝒞).

2. The property stated in Proposition 2.4 is not only a consequence of the fact that a subobject is characteristic, but it is equivalent to it, as shown in [13, Proposition 3.2]. Shortly:

H <_{char} G ⟺ (for each X, G ◁ X ⇒ H ◁ X) .

This is a consequence of Lemma 2.7 below.

2.7. Lemma. [13, Lemma 2.6] Consider a split extension as in the bottom row of the diagram

[figure: H; h; Y; B; G; k; X; p; B; s]

such that kh is normal. Then this split extension lifts along h: H → G to yield a normal monomorphism of split extensions.

Proof. The needed lifting is obtained via the pullback of split extensions in the diagram

[figure: H; h; Y; B; s; H; kh; R; r_{1}; ⟨r_{1},r_{2}⟩; X; G; k; k; X; p; ⟨sp,1_{X}⟩; B; s; s; X; ⟨0,1_{X}⟩; X × X; π_{1}; X; ⟨1_{X},1_{X}⟩]

where R is the equivalence relation on X associated with the normal subobject kh.

Thanks to Proposition 2.2 and Remark 2.6, in a semi-abelian category, we have three equivalent formulations of the property, for a subobject, of being characteristic:

1. H <_{char} G if every (internal) action on G restricts to an action on H;

2. H <_{char} G if, for every point with kernel G in a fibre Pt_{B}(𝒞), the inclusion of H in G

lifts to a monomorphism in Pt_{B}(𝒞);

3. H <_{char} G if, whenever G ◁ K, then H ◁ K.

We chose the first one as a definition, because it is a more natural generalisation of the already existing notion for groups and for Lie algebras. Notice, however, that the three formulations need not be equivalent outside the semi-abelian setting, since, in general, actions are not equivalent to points and normal subobjects do not coincide with clots (i.e. those closed under conjugation). The more exportable definition is probably the third, which makes sense in any category where a notion of normal subobject is defined. In that case, we should specify which kind of “normality” we are considering, since there are different notions of normal subobject, that coincide in the semi-abelian context (see, for example, [20] for a detailed account). These would give different, possibly non-equivalent, definitions of characteristic subobject. The study of the relationship between these notions in a more general context goes beyond the purposes of this paper and is material for a future work. Accordingly, from now on, unless otherwise specified, 𝒞 will be a semi-abelian category and one can think of normal subobjects simply as kernels.

When the category 𝒞 is not only semi-abelian, but also strongly protomodular [7], internal actions behave well with respect to quotients. More precisely, in [21] the following result is proved.

2.8. Proposition. [21, Theorem 5.3] A semi-abelian category is strongly semi-abelian (i.e. semi-abelian and strongly protomodular) if and only if the following property holds:

• for every normal subobject H ◁ G and every action ξ: B♭G → G, if ξ restricts to H, then ξ also induces a (unique) action ξ̃ on the quotient G/H:

[figure: B♭H; \overline{ξ}; B♭G; ξ; B♭(G/H); ξ̃; H; h; G; q; G/H]

In terms of split extensions, this means that if a kernel h is the restriction of some ϕ in Pt_{B}(𝒞), then q = coker(h) is the restriction of γ = coker(ϕ) in Pt_{B}(𝒞):

[figure: H; h; Y; ϕ; B; G; q; X; γ; B; G/H; Z; B]  (2)

It turns out that, for the special class of characteristic subobjects, strong protomodularity is not needed in order to transfer actions to the quotient.

2.9. Proposition. If H is a characteristic subobject of G, then every action on G induces an action on the quotient G/H, as in the diagram of Proposition 2.8.

Proof. As already observed in Remark 2.6, for every action of an object B on G, there is a normal monomorphism in Pt_{B}(𝒞) whose restriction to the kernels is h. By taking its cokernel, we get an exact sequence in Pt_{B}(𝒞) as in diagram (2).

2.10. Proposition. If H ≤ K ≤ G, H is characteristic in G and K/H is characteristic in G/H, then K is characteristic in G.

Proof. Let us consider the following diagram

[figure: H; K; k; K/H; k̃; H; G; q; G/H]

The right hand side square is a pullback (this comes from the fact that the category 𝒞, being semi-abelian, is protomodular [5]). By Proposition 2.9, every action of some B on G induces an action on G/H. By assumption, the same action restricts to K/H. In terms of points, we have a cospan in Pt_{B}(𝒞) whose restriction to the kernels is the pair (q, k̃). Now, since the kernel functors preserve pullbacks, K is the kernel of the pullback in Pt_{B}(𝒞) of the same cospan, hence the action of B on G restricts to K.

2.11. Proposition. If H is characteristic in G, then its corresponding equivalence relation R on G is closed under actions on G, i.e. there exists an action R(ξ) of B on R which makes the following diagram commute:

\xymatrix{B♭R \ar[d]_{R(ξ)} \ar[r]^{1♭r_{1}} \ar[r]_{1♭r_{2}} & B♭G \ar[d]^{ξ} \\ R \ar[r]^{r_{1}} \ar[r]_{r_{2}} & G}

Proof. By Proposition 2.9, every action of an object B on G induces an action on G/H. Now, since kernel functors preserve pullbacks, R is the kernel of the kernel pair in Pt_{B}(𝒞) of the morphism γ of diagram (2):

[figure: R; r_{1}; r_{2}; R_{γ}; B; G; q; X; γ; B; G/H; Z; B]

We can make explicit the previous proposition in the category Gp. It says that for all pairs (x, y) ∈ R and for all b ∈ B, the pair (^{b}x, ^{b}y) belongs to R.

More in general, whenever B acts on G, there is an induced action on G × G (simply computing the product in Pt_{B}(𝒞)), and the inclusion R ↣ G × G is compatible with the corresponding actions. However, this does not mean that R is a characteristic subobject of G × G.

2.12. Proposition. If H and K are characteristic subobjects of G, then their intersection H ∧ K is characteristic in G.

Proof. Thanks to Remark 2.6, this is an immediate consequence of the fact that the intersection of normal subobjects is normal.

The result above can be extended to infinite families, provided that the infinite intersection exists.

2.13. Proposition. If H and K are characteristic subobjects of G, then their join H ∨K is characteristic in G.

Proof. Again, via Remark 2.6, this is an immediate consequence of the fact that, in a semi-abelian category, the join of two normal subobjects is normal (see [1]).

3. Commutators

While the outcomes listed in Section 2 hold in the very general case of semi-abelian categories, other classical properties of characteristic subgroups can be extended only under additional requirements on the base category.

An assumption, which turns out to be crucial in this sense, is to ask that kernel functors preserve jointly strongly epimorphic pairs. This is equivalent to the fact that, for all pairs ((Y, p_{1}, s_{1}), (Z, p_{2}, s_{2})) of objects in Pt_{B}(𝒞), the canonical arrow in 𝒞:

Ker_{B}(Y, p_{1}, s_{1}) + Ker_{B}(Z, p_{2}, s_{2}) → Ker_{B}((Y, p_{1}, s_{1}) + (Z, p_{2}, s_{2}))

is a regular epimorphism. This condition, which was first considered in the present paper and in [15], has been recently called algebraic coherence in [12], where it is studied in details. An analogous condition, namely the preservation of jointly epimorphic pairs by the functor B♭− for any B, was already considered in [19].

A context in which the property of preservation of jointly strongly epimorphic pairs by the kernel functors holds is the one of locally algebraically cartesian closed categories [8]. A semi-abelian category 𝒞 is said locally algebraically cartesian closed (or simply LACC) if, for any morphism f: A → B in 𝒞, the change of base functor

f^{∗}: Pt_{B}(𝒞) → Pt_{A}(𝒞) ,

defined by taking pullbacks along f, has a right adjoint. Examples of this situation are the categories Gp of groups and R-Lie of Lie algebras over a fixed commutative ring R. In this context the kernel functors (which are change of base functors with A = 0), having right adjoints, preserve all finite colimits, and hence the canonical arrow

Ker_{B}(Y, p_{1}, s_{1}) + Ker_{B}(Z, p_{2}, s_{2}) → Ker_{B}((Y, p_{1}, s_{1}) + (Z, p_{2}, s_{2}))

mentioned above is an isomorphism.

Another context in which the kernel functors preserve jointly strongly epimorphic pairs is given by categories of interest [23], as proved in [12]. We recall that a category of interest in the sense of [23] is a category 𝒞 whose objects are groups with a set of operation Ω and with a set of equalities 𝔼, such that 𝔼 includes the group laws and the following conditions hold. If Ω_{i} is the set of i-ary operations in Ω, then:

(a) Ω = Ω_{0} ∪ Ω_{1} ∪ Ω_{2};

(b) the group operations (written additively: 0, −, +, even if the group is not necessarily abelian) are elements of Ω_{0}, Ω_{1} and Ω_{2} respectively. Let Ω′_{2} = Ω_{2}\{+}, Ω′_{1} = Ω_{1}\{−} and assume that if ∗ ∈ Ω′_{2}, then Ω′_{2} contains ∗^{∘} defined by x ∗^{∘} y = y ∗ x. Assume further that Ω_{0} = {0};

(c) for any ∗ ∈ Ω′_{2}, 𝔼 includes the identity x ∗ (y + z) = x ∗ y + x ∗ z;

(d) for any ω ∈ Ω′_{1} and ∗ ∈ Ω′_{2}, 𝔼 includes the identities ω(x + y) = ω(x) + ω(y) and ω(x) ∗ y = ω(x ∗ y);

(e) Axiom 1 x_{1} + (x_{2} ∗ x_{3}) = (x_{2} ∗ x_{3}) + x_{1} for any ∗ ∈ Ω′_{2};

(f) Axiom 2 for any ordered pair (∗, \overline{∗}) ∈ Ω′_{2} × Ω′_{2} there is a word W such that

(x_{1} ∗ x_{2})\overline{∗}x_{3} = W (x_{1}(x_{2}x_{3}), x_{1}(x_{3}x_{2}), (x_{2}x_{3})x_{1}, (x_{3}x_{2})x_{1},
x_{2}(x_{1}x_{3}), x_{2}(x_{3}x_{1}), (x_{1}x_{3})x_{2}, (x_{3}x_{1})x_{2}),

where each juxtaposition represents an operation in Ω′_{2}.

Examples of categories of interest are groups, Lie algebras, rings, associative algebras, Leibniz algebras, Poisson algebras and many others.

Since it will be useful later, we give here a description of internal actions in this context (called derived actions in [11]). In a category of interest 𝒞, an action of an object B on an object X is a set of functions:

f_{∗}: B × X → X ,

one for each operation ∗ in Ω_{2} (we will write b · x for f_{+}(b, x) and b ∗ x for f_{∗}(b, x), with ∗ ∈ Ω′_{2}), such that the one corresponding to the group operation + satisfies the usual axioms for a group action, the others are bilinear with respect to + and moreover the following axioms are satisfied (for all b, b_{i} ∈ B, x, x_{i} ∈ X and ∗, \overline{∗} ∈ Ω′_{2}):

1. b · (x_{1} ∗ x_{2}) = x_{1} ∗ x_{2};

2. x_{1} + (b ∗ x_{2}) = (b ∗ x_{2}) + x_{1};

3. (b_{1} ∗ b_{2}) · x = x;

4. b_{1} · (b_{2} ∗ x) = b_{2} ∗ x;

5. (b ∗ x_{1})\overline{∗}x_{2} = W (b(x_{1}x_{2}), b(x_{2}x_{1}), (x_{1}x_{2})b, (x_{2}x_{1})b, x_{1}(bx_{2}), x_{1}(x_{2}b), (bx_{2})x_{1}, (x_{2}b)x_{1});

6. (x_{1} ∗ x_{2})\overline{∗}b = W (x_{1}(x_{2}b), x_{1}(bx_{2}), (x_{2}b)x_{1}, (bx_{2})x_{1}, x_{2}(x_{1}b), x_{2}(bx_{1}), (x_{1}b)x_{2}, (bx_{1})x_{2});

7. (b_{1} ∗ b_{2})\overline{∗}x = W (b_{1}(b_{2}x), b_{1}(xb_{2}), (b_{2}x)b_{1}, (xb_{2})b_{1}, b_{2}(b_{1}x), b_{2}(xb_{1}), (b_{1}x)b_{2}, (xb_{1})b_{2});

8. (b_{1} ∗ x)\overline{∗}b_{2} = W (b_{1}(xb_{2}), b_{1}(b_{2}x), (xb_{2})b_{1}, (b_{2}x)b_{1}, x(b_{1}b_{2}), x(b_{2}b_{1}), (b_{1}b_{2})x, (b_{2}b_{1})x);

where W indicates the same word in Axiom 2 corresponding to the choice of ∗ and \overline{∗}.

Observe that axioms 1–4 above come from Axiom 1, while axioms 5–8 come from Axiom 2 by replacing each operation with the corresponding action (notice that the group action replaces the conjugation and not the group operation). These axioms are nothing but the translation of the condition that one obtains by considering the equivalence between actions and points and expressing the action as the conjugation into the semidirect product. More explicitly, given a split extension:

\xymatrix{X \ar[r]^{k} & A \ar[r]^{p} & B , \ar[l]^{s}}

the corresponding action is given by:

b · x = k^{−1}(s(b) + k(x) − s(b));
b ∗ x = k^{−1}(s(b) ∗ k(x)).

A wider class of semi-abelian varieties is given by groups with operations introduced by Porter in [24]. In that class, the description of internal actions is similar to the one given above; axioms 1–8 are replaced by suitable ones coming from the identities of the corresponding algebraic theory. We will use this description explicitly in the Examples 3.4 and 4.5 below.

A classical property of characteristic subgroups of a group is the fact that the commutator of two characteristic subgroups is characteristic as well. In order to study this property in a semi-abelian setting, we will use an intrinsic definition of commutator of two subobjects. There are different possible definitions. The first we consider is the Huq commutator [17], which can be described in the following way (see [1] and [20]): given two subobjects h: H ↣ G and k: K ↣ G of an object G, we can construct the following diagram:

[figure: H + K; Σ_{H,K}; [h,k]; H × K; [H, K]_{G}; G; π; \frac{G}{[H,K]_{G}}]

where Σ_{H,K} is the canonical map

Σ_{H,K} = ⟨[1, 0], [0, 1]⟩ = [⟨1, 0⟩, ⟨0, 1⟩]: H + K → H × K

and the commutative square is a pushout. Then the Huq commutator appears as the kernel of the morphism π. Being a kernel, the Huq commutator is always a normal subobject, even if H and K are not.

Another possible definition is that of Higgins commutator [20]. Given two subobjects h: H ↣ G and k: K ↣ G of an object G, let us denote by σ_{H,K}: H ⋄ K → H + K the kernel of the canonical morphism Σ_{H,K}: H + K → H × K. The Higgins commutator [H, K] of H and K is the regular image of H ⋄ K under the morphism [h, k]σ_{H,K}, as in the following diagram:

[figure: H ⋄ K; σ_{H,K}; H + K; [h,k]; [H, K]; G]

The Higgins commutator of H and K is not necessarily a normal subobject of G, even when H and K are. In fact, its normal closure in G is the Huq commutator. Following [13], we say that a category 𝒞 satisfies the (NH) property when the Higgins commutator of two normal subobjects is normal, or, in other words, when Higgins and Huq commutators of normal subobjects coincide. The (NH) property is fulfilled both by (LACC) categories and by categories of interest (see [13]).

Let us observe that, if G coincides with the join of its subobjects H and K (or, in other words, when [h, k] is a regular epimorphism), then the Higgins commutator [H, K] is normal in G, since in the semi-abelian context regular images of normal subobjects along regular epimorphisms are normal. As a special case, this happens when h = k = 1_{G}, showing that the derived object [G, G] is always normal in G.

3.1. Proposition. Let 𝒞 be a semi-abelian category where the kernel functors preserve jointly strongly epimorphic pairs. If H and K are characteristic subobjects of G, then the Higgins commutator [H, K] is a characteristic subobject of G. In particular, the derived subobject [G, G] is a characteristic subobject of G.

Proof. If H and K are characteristic subobjects of G, then, for every action ξ: B♭G → G, there is a cospan in Pt_{B}(𝒞):

[figure: H; k_{1}; h; Y; p_{1}; B; s_{1}; G; i_{G}; G ⋊_{ξ} B; p_{B}; B; i_{B}; K; k_{2}; k; Z; p_{2}; B; s_{2}]

The product (Y, p_{1}, s_{1}) × (Z, p_{2}, s_{2}) in Pt_{B}(𝒞) has H × K as kernel. In general, the kernel N of the coproduct (Y, p_{1}, s_{1}) + (Z, p_{2}, s_{2}) is different from H + K; however, under our hypothesis, the canonical map u: H+K → N is a regular epimorphism. Now, consider the following commutative diagram, where α is the arrow induced on kernels by the canonical morphism (Y, p_{1}, s_{1}) + (Z, p_{2}, s_{2}) → (Y, p_{1}, s_{1}) × (Z, p_{2}, s_{2}) in Pt_{B}(𝒞), β is induced by (Y, p_{1}, s_{1}) + (Z, p_{2}, s_{2}) → (G ⋊ B, p_{B}, i_{B}), and j = ker(α):

[figure: H ⋄ K; σ_{H,K}; v; H + K; Σ_{H,K}; ⌜; u; H × K; M; j; N; α; β; H × K; [H, K]; G]

The arrow v: H ⋄ K → M is a regular epimorphism, since the square αu = 1_{H×K}Σ_{H,K} is a pushout. The Higgins commutator [H, K] in G is defined as the regular image of σ_{H,K} along βu, which is the same as the regular image of j along β. Now recall that the composite βj is the restriction to kernels of a morphism in Pt_{B}(𝒞). Since the (regular epi, mono) factorisation of a morphism in Pt_{B}(𝒞) is preserved by the kernel functor, then the inclusion [H, K] ↣ G turns out to be the restriction to kernels of a morphism in Pt_{B}(𝒞), thus being characteristic as follows from Proposition 2.2.

3.2. Corollary. Let 𝒞 be a semi-abelian category where the kernel functors preserve jointly strongly epimorphic pairs. If H and K are characteristic subobjects of an object G in 𝒞 then the Huq commutator and the Higgins commutator of H and K in G coincide.

Proof. This is a consequence of Proposition 3.1 and Corollary 2.5, since the Huq commutator is the normal closure of the Higgins commutator.

The results above apply, in particular, either to a semi-abelian (LACC) category or to a category of interest. This depends on the fact that both classes of categories satisfy the hypothesis of Proposition 3.1, as explained above.

For example, in the category of (not necessarily unitary) rings, given a ring X and two subrings H and K, the commutator [H, K] is nothing but the subring HK of X generated by all elements of the form hk or kh, with h ∈ H and k ∈ K. Hence Proposition 3.1 says that, if H and K are characteristic, HK also is. Something similar happens in the category of Lie algebras (over a commutative ring R), where the commutator [H, K] of two subalgebras is the Lie subalgebra generated by all elements of the form [h, k], with h ∈ H and k ∈ K.

3.3. Remark. In fact, in a semi-abelian context, the property:

H, K characteristic in X ⇒ [H, K] characteristic in X

is also implied by the (NH) property, as shown in [13, Proposition 3.3]. This gives an alternative proof in the case of (LACC) categories and of categories of interest.

The fact that the Higgins (or the Huq) commutator of two characteristic subobjects is characteristic is not true in a general semi-abelian category. Not even the derived subobject of an object (which is the same in the Higgins or in the Huq sense) is characteristic in general, as the following example shows. On the other hand, Example 3.5 below shows that, even in the category of groups, the commutator [H, K] fails to be characteristic if H and K are not characteristics.

3.4. Example. Let us consider the category NARng of not necessarily associative rings, i.e. abelian groups with a binary operation which is distributive over the group operation. Let us consider the object G in NARng given by the free abelian group on two generators G = ℤx + ℤy, endowed with a distributive binary operation, defined on generators as:

\begin{matrix}∗ & x & y \\ x & x & 0 \\ y & 0 & 0 \end{matrix}

Then the derived subobject [G, G] = ℤx is an ideal (i.e. a normal subobject) of G, but it is not characteristic in G. Indeed, if we consider the object given by the abelian group ℤ with trivial multiplication, [G, G] is not stable under the following action of ℤ over G:

\begin{matrix}ℤ × G → G , & z ∗ (αx + βy) = (zβ)x + (zα)y , \\ G × ℤ → G , & (αx + βy) ∗ z = (zβ)x + (zα)y . \end{matrix}

We emphasize that G is, in fact, an associative ring, but the present is not a counterexample for the category Rng of rings, since the one described above is an action in NARng but not in Rng. Indeed, according to the explicit description of actions recalled at the beginning of this section, an action of ℤ over G in NARng is just a pair of bilinear maps ℤ×G → G and G×ℤ → G, while an action in Rng must also satisfy some “associativity” axioms. In the example above, the axiom

z ∗ (xx) = (z ∗ x)x

is not satisfied, indeed z ∗ (xx) = z ∗ x = zy, while (z ∗ x)x = (zy)x = 0.

3.5. Example. Let S_{3} be the symmetric group over the set of three objects, h: A_{3} → S_{3} its normal subgroup of even permutation. Let G = S_{3} × S_{3} and H be the subgroup A_{3} with the inclusion ⟨h, 0⟩, which is normal. One can prove that [G, H] = H, which is not a characteristic subobject of G, since it is not fixed by the twisting automorphism ⟨π_{2}, π_{1}⟩: S_{3} × S_{3} → S_{3} × S_{3}.

4. Centres and centralisers

Given a characteristic subgroup H of a group G, its centraliser Z_{G}(H) is characteristic, too. In particular, the centre of a group is always a characteristic subgroup. This is not true in any semi-abelian category, as we will show later, so we need to consider further hypotheses on the category in order to get this property. In a semi-abelian category 𝒞, given a subobject H of an object G, the centraliser of H in G is the largest subobject Z_{G}(H) of G such that the Huq commutator [H, Z_{G}(H)]_{G} vanishes. The centre of an object G is the largest subobject Z(G) of G such that [G, Z(G)] = 0.

Centres and centralisers do not always exist in a semi-abelian category, and even when they exist, they can be difficult to handle. Bourn and Janelidze introduced in [10] a categorical context, namely action accessible categories, in which the centres and the centralisers have an easy description. We recall now the definition of action accessible categories and their basic properties.

Let 𝒞 be a semi-abelian category. Fixed an object K ∈ 𝒞, a split extension with kernel K is a diagram

\xymatrix{K \ar[r]^{k} & A \ar[r]^{p} & B \ar[l]^{s}} ,

such that ps = 1_{B} and k = Ker(p). We denote such a split extension by (B, A, p, s, k). Given another split extension (D, C, q, t, l) with the same kernel K, a morphism of split extensions

(g, f): (B, A, p, s, k) ⟶ (D, C, q, t, l)  (3)

is a pair (g, f) of morphisms:

[figure: K; k; A; f; p; B; g; s; K; l; C; q; D; t]  (4)

such that l = fk, qf = gp and fs = tg. Let us notice that, since the category 𝒞 is protomodular, the pair (k, s) is jointly (strongly) epimorphic, and then the morphism f in (4) is uniquely determined by g.

The split extensions with fixed kernel K form a category, denoted by SplExt_{𝒞}(K), or simply by SplExt(K).

4.1. Definition. [10]

• An object in SplExt(K) is said to be faithful if any object in SplExt(K) admits at most one morphism into it.

• Split extensions with a morphism into a faithful one are called accessible.

• If, for any K ∈ 𝒞, every object in SplExt(K) is accessible, we say that the category 𝒞 is action accessible.

In the case of groups, faithful extensions are those inducing a group action of B on K (via conjugation in A) which is faithful. Every split extension in Gp is accessible and a morphism into a faithful one can be performed by taking the quotients of B and A over the centraliser Z_{B}(K), i.e. the (normal) subobject of A given by those elements of B that commute in A with every element of K.

In [22] it is proved that any category of interest in the sense of [23] is action accessible. Examples of action accessible categories are then groups, rings, associative algebras, Lie algebras, Leibniz algebras and Poisson algebras, as mentioned before.

In the context of action accessible categories it is easy to describe the centraliser of a normal subobject. We give now a brief description of the construction, without proof (that can be found, for example, in [14]). Let x: X → A be a normal subobject of A, and let R[p] be the equivalence relation on A induced by X (i.e. the kernel pair of the quotient p: A → A/X). Consider the following morphism of split extensions, where the codomain is a faithful one (it exists because the category is action accessible):

[figure: X; ⟨x,0⟩; R[p]; r_{0}; f; A; s_{0}; g; X; k; C; q; D; t]

Then the kernel of g is the centraliser Z_{A}(X) of X in A. This implies, in particular, that in an action accessible category the centraliser of a normal subobject is normal [14, Corollary 2.6], which is not always the case in general semi-abelian categories, even when Z_{A}(X) exists (see examples in [14]).

We are now ready to prove that, in the context of action accessible categories, the centraliser of a characteristic subobject is characteristic.

4.2. Lemma. Let 𝒞 be a semi-abelian category where, for every normal subobject H ◁ G, the centraliser Z_{G}(H) of H in G is normal in G. Then if G′ is a normal subobject of G, Z_{G′}(H) is also normal in G.

Proof. By definition of centraliser, Z_{G′}(H) is the largest subobject of G′ such that [H, Z_{G′}(H)]_{G′} = 0. Hence, it is contained in both G′ and Z_{G}(H), and it is the largest with this property, so it is defined by the following pullback:

[figure: Z_{G′}(H); Z_{G}(H); G′; G]

In other words, Z_{G′}(H) = Z_{G}(H) ∧ G′ and it is normal in G as intersection of two normal subobjects.

4.3. Proposition. Let 𝒞 be a semi-abelian category where, for every normal subobject H ◁ G, the centraliser Z_{G}(H) of H in G is normal in G. Then if H is a characteristic subobject of G, Z_{G}(H) is also characteristic in G.

Proof. Consider an object B and an action ξ: B♭G → G. G is a normal subobject of G ⋊_{ξ} B; so, being characteristic in G, H is normal in G ⋊_{ξ} B by Proposition 2.4. Hence, by Lemma 4.2, Z_{G}(H) is a normal subobject of G ⋊_{ξ} B. Now, we can apply Lemma 2.7 to the following situation:

[figure: Z_{G}(H); 0; G; i_{G}; G ⋊_{ξ} B; p_{B}; B; i_{B}; 0]

thus obtaining a morphism of split extensions:

[figure: 0; Z_{G}(H); Z_{G⋊_{ξ}B}(G); B; 0; 0; G; i_{G}; G ⋊_{ξ} B; p_{B}; B; i_{B}; 0]

which gives the desired action ξ′: B♭Z_{G}(H) → Z_{G}(H) as a restriction of the action ξ.

4.4. Corollary. Let 𝒞 be a semi-abelian category where, for every normal subobject H ◁ G, the centraliser Z_{G}(H) of H in G is normal in G. Then the centre Z(G) is a characteristic subobject of G.

In the category of (not necessarily unitary) rings, given an ideal H of a ring G, the centraliser Z_{G}(H) is the annihilator of H in G, i.e.

Z_{G}(H) = {g ∈ G | gh = hg = 0 for all h ∈ H} .

Hence, if H is characteristic in G, then the annihilator of H in G is characteristic, as well. In particular, for any ring G, the annihilator of G is a characteristic ideal of G. The same happens in the category of Lie algebras over a commutative ring R.

Proposition 4.3 and Corollary 4.4 are true, in particular, in semi-abelian action accessible categories. However, they do not hold in any semi-abelian category. The following is a counterexample.

4.5. Example. Let us consider again the category NARng of not necessarily associative rings and the object G in NARng described in Example 3.4. The centre Z(G) = ℤy is an ideal (i.e. a normal subobject) of G, but it is not characteristic in G, since it is not stable under the action of ℤ over G described in the same example.

5. Induced morphisms between actors

In the category Gp of groups, if H is a characteristic subgroup of G, then there are induced morphisms Aut(G) → Aut(H) and Aut(G) → Aut(G/H). This comes from the fact that actions on G (which are equivalent to split extensions with kernel G, as already observed) are represented by the automorphism group Aut(G), in the sense that an action of a group B on G can be described simply as a group homomorphism B → Aut(G). We are going to show that the same induced morphisms exist in a context in which internal actions are representable in the sense of [3, 4]. Categories in which this happens are called action representative in [2]. Let us recall the definition.

5.1. Definition. [3, 2] A semi-abelian category 𝒞 is action representative if, for any object X ∈ 𝒞, there exists an object Act(X), called the actor of X, and a split extension

\xymatrix{X \ar[r] & X ⋊ Act(X) \ar[r] & Act(X) \ar[l]} ,

called the split extension classifier of X, which is terminal in SplExt(X). That is, for any split extension with kernel X:

\xymatrix{X \ar[r]^{k} & A \ar[r]^{p} & B \ar[l]^{s}}

there exists a unique morphism (φ, φ_{1}) of split extensions from (B, A, p, s, k) to the split extension classifier:

[figure: X; k; A; φ_{1}; p; B; s; φ; X; X ⋊ Act(X); Act(X)]

Notice that the morphism φ_{1} is uniquely determined by φ and the identity on X (since k and s are jointly strongly epimorphic).

Examples of action representative categories are the category Gp of groups, where the actor is the group of automorphisms, and the category Lie of Lie algebras over a commutative ring R, where the actor of a Lie algebra X is the Lie algebra Der(X) of derivations of X.

It is well-known that the assignment G ↦ Act(G) is not functorial. Nevertheless, it behaves well with respect to characteristic subobjects.

5.2. Proposition. Let 𝒞 be an action representative semi-abelian category. Every characteristic subobject h: H ↣ G induces a morphism between split extension classifiers:

[figure: G; q; G ⋊ Act(G); Act(G); G/H; G/H ⋊ Act(G/H); Act(G/H)]  (5)

and a morphism between actors: Act(G) → Act(H).

Proof. As explained in Section 2, if H is a characteristic subobject of G, then, for every action ξ: B♭G → G, there exists an exact sequence in Pt_{B}(𝒞):

[figure: H; h; Y; ϕ; B; G; q; X; γ; B; G/H; Z; B]

Since the category 𝒞 is action representative, we can choose, in particular, B = Act(G) and the middle row to be the split extension classifier of G. Thus, thanks to Proposition 2.9, we have a morphism in Pt_{Act(G)}(𝒞):

[figure: G; G ⋊ Act(G); Act(G); G/H; Z; Act(G)]

By composing with the arrow to the split extension classifier of G/H, we get the desired morphism (5).

For the same reason, we also have a morphism:

[figure: H; h; Y; Act(G); G; G ⋊ Act(G); Act(G)]

The arrow from the upper split extension to the split extension classifier of H produces the morphism Act(G) → Act(H).

It is worth translating the above proposition in terms of internal actions. The first assertion says that there exists a morphism q̃: Act(G) → Act(G/H) making the following diagram commute:

\xymatrix{Act(G)♭G \ar[r]^{q̃♭q} \ar[d]_{ζ_{G}} & Act(G/H)♭(G/H) \ar[d]^{ζ_{G/H}} \\ G \ar@{|->}[r]_{q} & G/H}

where ζ_{G} and ζ_{G/H} are the canonical actions of the actors. On the other hand, the second statement says that there exists a morphism \overline{h}: Act(G) → Act(H) making this triangle commute:

\xymatrix{Act(G)♭H \ar[r]^{\overline{h}♭1} \ar[dr]_{\overline{ζ_{G}}} & Act(H)♭H \ar[d]^{ζ_{H}} \\  & H}

where \overline{ζ_{G}} is the action on H induced by ζ_{G} and ζ_{H} is the canonical action of the actor.

Let us observe that any action representative category is action accessible: indeed, it is easy to see that the split extension classifier is a faithful split extension. On the other hand, the category Rng of rings is action accessible [10] but not action representative. In the case of action accessible categories, one cannot recover the same properties described above for action representative categories, because there can be many faithful split extensions associated with a given one. However, as observed in [14], there always exists a canonical faithful split extension associated with a given one, and it has properties analogous to the ones described above.

In an action accessible semi-abelian category, given a morphism of split extensions with faithful codomain:

[figure: X; k; A; f; p; B; s; g; X; C; q; D; t]

the canonical (regular epi, mono) factorisation gives rise to another faithful split extension:

[figure: X; k; A; e_{f}; p; B; s; e_{g}; X; T_{1}; m_{f}; T_{0}; m_{g}; X; C; q; D.; t]

The important fact here is that the faithful split extension in the middle of the previous diagram does not depend on the choice of the lower one, so it is a canonical faithful split extension associated with (A, B, p, s). The object T_{0} is actually the quotient B/Z_{B}(X) of B over the centraliser of X in B (i.e. the largest subobject of B commuting with X in A), while T_{1} is the quotient A/Z_{B}(X).

As above, let H be a characteristic subobject of G. Then, for every action ξ of B on G, there exists an exact sequence in Pt_{B}(𝒞) as in diagram (2). Let

[figure: G; X; B; G; T_{1}(B, G, ξ); T_{0}(B, G, ξ)]

be the morphism onto the canonical faithful split extension (and similarly for the induced split extensions of kernels H and G/H).

5.3. Proposition. Let 𝒞 be an action accessible semi-abelian category. Every characteristic subobject h: H ↣ G induces a morphism between canonical faithful split extensions:

[figure: G; q; T_{1}(B, G, ξ); T_{0}(B, G, ξ); G/H; T_{1}(B, G/H, ξ̃); T_{0}(B, G/H, ξ̃)]  (6)

and a morphism: T_{0}(B, G, ξ) → T_{0}(B, H, \overline{ξ}).

Proof. As explained above, the object T_{0}(B, G, ξ) is nothing but the quotient B/Z_{B}(G), and T_{1}(B, G, ξ) ≅ X/Z_{B}(G), and similarly for T_{i}(B, H, \overline{ξ}) and T_{i}(B, G/H, ξ̃). The desired morphism (6) will be the bottom rectangle in the following commutative diagram:

[figure: G; q; X; γ; B; G/H; Z; B; G; q; T_{1}(B, G, ξ); q_{1}; T_{0}(B, G, ξ); q_{0}; G/H; T_{1}(B, G/H, ξ̃); T_{0}(B, G/H, ξ̃)]

It is constructed as follows. By definition, the centraliser Z_{B}(G) of G in B is such that [G, Z_{B}(G)]_{X} = 0. Composing with γ, we also have [G/H, Z_{B}(G)]_{Z} = 0, so that Z_{B}(G) ≤ Z_{B}(G/H), and this induces the arrow q_{0} between the corresponding cokernels over B. On the other hand, q_{1} is the arrow which completes the following morphism of

226  ALAN S. CIGOLI AND ANDREA MONTOLI

short exact sequences:

[figure: Z_{B}(G); X; γ; T_{1}(B, G, ξ); q_{1}; Z_{B}(G/H); Z; T_{1}(B, G/H, ξ̃)]

To prove the second assertion, consider the morphism below in Pt_{B}(𝒞):

[figure: H; h; Y; ϕ; B; G; X; B]

By definition, [G, Z_{G}(B)]_{X} = 0 and, as a consequence, [H, Z_{G}(B)]_{X} = 0. Since ϕ is monomorphic, this implies that [H, Z_{G}(B)]_{Y} = 0 too, hence Z_{G}(B) ≤ Z_{H}(B). The morphism T_{0}(B, G, ξ) → T_{0}(B, H, \overline{ξ}) is the one induced on the corresponding cokernels over B.

6. Summarising table

Property  True in  Reference
H <_{char} G ⇒ H ◁ G  𝒞 semi-abelian  2.5
H <_{char} K ◁ G ⇒ H ◁ G  𝒞 semi-abelian  2.4
H <_{char} K <_{char} G ⇒ H <_{char} G  𝒞 semi-abelian  2.3
H, K <_{char} G ⇒ H ∧ K <_{char} G  𝒞 semi-abelian  2.12
H, K <_{char} G ⇒ H ∨ K <_{char} G  𝒞 semi-abelian  2.13
H <_{char} G, B acts on G ⇒ B acts on G/H  𝒞 semi-abelian  2.9

H ≤ K ≤ G, H <_{char} G, K/H <_{char} G/H

⇒ K <_{char} G

𝒞 semi-abelian 2.10

H <_{char} G, (R kernel pair of G → G/H)

⇒ R closed under actions on G

𝒞 semi-abelian 2.11

[G, G] <_{char} G

𝒞 semi-abelian algebraically coherent

3.1

H, K <_{char} G ⇒ [H, K] <_{char} G

𝒞 semi-abelian algebraically coherent

3.1

Z(G) <_{char} G 𝒞 semi-abelian action accessible

4.4

H <_{char} G ⇒ Z_{G}(H) <_{char} G 𝒞 semi-abelian action accessible

4.3

H <_{char} G ⇒ \{ \begin{matrix}Act(G) → Act(G/H) \\ Act(G) → Act(H) \end{matrix}

𝒞 semi-abelian action representative 5.2

References

[1] F. Borceux, D. Bourn, Mal’cev, protomodular, homological and semi-abelian categories, Mathematics and its applications, vol. 566 (2004), Kluwer.

[2] F. Borceux, D. Bourn, Split extension classifier and centrality, Contemp. Math., vol. 431 (2007) 85–104, Amer. Math. Soc.

[3] F. Borceux, G. Janelidze, G. M. Kelly, Internal object actions, Comment. Math. Univ. Carolin. 46 (2005) 235–255.

[4] F. Borceux, G. Janelidze, G. M. Kelly, On the representability of actions in a semiabelian category, Theory Appl. Categ. 14 (2005) 244–286.

[5] D. Bourn, Normalization equivalence, kernel equivalence and affine categories, Lecture Notes in Mathematics, vol. 1488 (1991) 43–62, Springer-Verlag.

[6] D. Bourn, Normal subobjects and abelian objects in protomodular categories, J. Algebra 228 (2000) 143–164.

[7] D. Bourn, Normal functors and strong protomodularity, Theory Appl. Categ. 7 (2000) 206–218.

[8] D. Bourn, J. R. A. Gray, Aspects of algebraic exponentiation, Bull. Belg. Math. Soc. Simon Stevin 19 (2012) 823–846.

[9] D. Bourn, G. Janelidze, Protomodularity, descent, and semidirect products, Theory Appl. Categ. 4 (1998) 37–46.

[10] D. Bourn, G. Janelidze, Centralizers in action accessible categories, Cah. Topol. Géom. Différ. Catég. 50 (2009) 211–232.

[11] J. M. Casas, T. Datuashvili, M. Ladra, Universal strict general actors and actors in categories of interest, Appl. Categ. Structures 18 (2010) 85–114.

[12] A. S. Cigoli, J. R. A. Gray, T. Van der Linden, Algebraically coherent categories, (2014), arXiv:1409.4219.

[13] A. S. Cigoli, J. R. A. Gray, T. Van der Linden, On the normality of Higgins commutators, J. Pure Appl. Algebra 219 (2015) 897–912.

[14] A. S. Cigoli, S. Mantovani, Action accessibility via centralizers, J. Pure Appl. Algebra 216 (2012) 1852–1865.

[15] A. S. Cigoli, S. Mantovani, G. Metere, Peiffer product and Peiffer commutator for internal precrossed modules, (2014), in preparation.

[16] J. R. A. Gray, Algebraic exponentiation in general categories, Appl. Categ. Structures 20 (2012) 543–567.

[17] S. A. Huq, Commutator, nilpotency and solvability in categories, Quart. J. Math. Oxford 19 (1968) 363–389.

[18] G. Janelidze, L. Márki, W. Tholen, Semi-abelian categories, J. Pure Appl. Algebra 168 (2002) 367–386.

[19] S. Mantovani, G. Metere, Internal crossed modules and Peiffer condition, Theory Appl. Categ. 23 (2010) 113–135.

[20] S. Mantovani, G. Metere, Normalities and commutators, J. Algebra 324 (2010) 2568– 2588.

[21] G. Metere, A note on strong protomodularity, actions and quotients, (2013), arXiv:1306.3773v1.

[22] A. Montoli, Action accessibility for categories of interest, Theory Appl. Categ. 23 (2010) 7–21.

[23] G. Orzech, Obstruction theory in algebraic categories, I, J. Pure Appl. Algebra 2 (1972) 287–314.

[24] T. Porter, Extensions, crossed modules and internal categories in categories of groups with operations, Proc. Edinburgh Math. Soc. 30 (1987) 373–381.

Dipartimento di Matematica, Università degli Studi di Milano Via Saldini 50, Milano, Italy

CMUC, Department of Mathematics, University of Coimbra 3001-501 Coimbra, Portugal

Email: alan.cigoli@unimi.it
      montoli@mat.uc.pt
This article may be accessed at http://www.tac.mta.ca/tac/

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors The typesetting language of the journal is TEX, and L^{A}TEX2e strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Ecole Polytechnique Fédérale de Lausanne:

gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis: cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown(at)btinternet.com

Valeria de Paiva: valeria.depaiva@gmail.com

Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Kathryn Hess, Ecole Polytechnique Fédérale de Lausanne: kathryn.hess@epfl.ch Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, Macquarie University: steve.lack@mq.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@buffalo.edu Tom Leinster, University of Edinburgh: Tom.Leinster@ed.ac.uk Ieke Moerdijk, Radboud University Nijmegen: i.moerdijk@math.ru.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Giuseppe Rosolini, Università di Genova: rosolini@disi.unige.it Alex Simpson, University of Edinburgh: Alex.Simpson@ed.ac.uk James Stasheff, University of North Carolina: jds@math.upenn.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: rfcwalters@gmail.com R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca